OpenStudy (anonymous):

What is the first step in solving 4(x + 2) = 10 + x?

OpenStudy (anonymous):

Distribute 4 into (x+2) to get 4x + 8.

OpenStudy (anonymous):

Distribute: 4(x + 2) 4x + 8 = 10 + x then subtract x from both sides 3x + 8 = 10 then subtract 8 from both sides 3x = 2 then divide both sides by 3 x = 2/3